# Quasi-Static IRS: 3D Shaped Beamforming for Area Coverage Enhancement

### v1

@@ II Phase-Tuning Mechanism and Design of QS-IRS
Figure 1: (a) RA manufactured by TMYTEK comprising 51$\times$51 static elements with different patch shapes/sizes[17]; (b) RA with hybrid polarization-phase tuning based on element mirroring/rotation[18]; (c) Electromechanical IRS with rotatable patches[19]; (d) Electromechanical IRS with movable element ground[20].
@@ III System Model
Consider far-field communications between the BS and a target area on the ground with severely blocked direct links. We focus on practical scenarios where a QS-IRS is appropriately deployed [^3] The case with multiple distributed IRSs[22] is left for future investigation. such that there exists LoS paths from the BS to the target area via the QS-IRS, as in [2],[4]. [^4] The non-LoS scenario with non-negligible direct links will be considered in our future work, whereby statistical analysis of link power and outage probability under (e.g., Rician) fading channel can be incorporated in our system model and optimization framework. Consider one omnidirectional antenna at the receiver, and one (synthesized) directional antenna (beam) at the BS with a peak power gain of $G_{\text{t}}$ pointing towards the QS-IRS. [^5] Our proposed method can be fully protocol-transparent as in [6], under given BS antenna configuration. Assume that $M\triangleq M_{\text{y}}\times M_{\text{z}}$ IRS elements are arranged as a uniform planar array (UPA) with $M_{\text{y}}$ and $M_{\text{z}}$ elements along the $-y\text{-}$ and $-z\text{-}$axis, respectively, and with the first element at the origin, as shown in Fig. 2. Consider the IRS ERP model [10] given by $GF(\Pi)$, where $G$ denotes the peak power gain and $F(\Pi)$ denotes the normalized power pattern from/to the azimuth and elevation angles $\Pi\triangleq(\varphi,\theta)$, i.e.,
@@ III-A Array Steering Vector
Denote $\boldsymbol{u}_{\text{i}}\triangleq[\cos\varphi_{\text{i}}\sin\theta_{\text{i}},\sin\varphi_{\text{i}}\sin\theta_{\text{i}},\cos\theta_{\text{i}}]^{T}$ as the unit direction vector of the BS position w.r.t. the origin, with $\Pi_{\text{i}}\triangleq(\varphi_{\text{i}},\theta_{\text{i}})$ denoting the corresponding (horizontal, vertical) angle. The components of the steering vector incident on the QS-IRS along the $-y\text{-}$ and $-z\text{-}$axis can be expressed as
@@
where $\boldsymbol{r}_{m_{\text{y}}}\triangleq[0,(m_{\text{y}}-1)d_{\text{y}},0]^{T}$ and $\boldsymbol{r}_{m_{\text{z}}}\triangleq[0,0,(m_{\text{z}}-1)d_{\text{y}}]^{T}$ denote the position vectors of the $(m_{\text{y}},m_{\text{z}})$-th element along the $-y\text{-}$ and $-z\text{-}$axis, with $d_{\text{y}}$ and $d_{\text{z}}$ being the element spacing; $c$ denotes the speed of light and $f_{\text{c}}$ denotes the carrier frequency. Then, the incident steering vector is given by $\boldsymbol{a}_{\text{i}}(\Pi_{\text{i}})\triangleq\boldsymbol{a}_{\text{i}\text{y}}(\Pi_{\text{i}})\otimes\boldsymbol{a}_{\text{i}\text{z}}(\Pi_{\text{i}})$. Similarly, the components of the reflect steering vector are given by
@@ III-B Channel Power Gain
where (11) is due to (III-A) and the decomposition of $\boldsymbol{w}=\boldsymbol{w}_{\text{y}}\otimes\boldsymbol{w}_{\text{z}}$, with $-y\text{-}$ component $\boldsymbol{w}_{\text{y}}\triangleq\left[e^{j\phi_{1}},\cdots,e^{j\phi_{M_{\text{y}}}}\right]^{T}$ and $-z\text{-}$ component $\boldsymbol{w}_{\text{z}}\triangleq\left[e^{j\phi_{1}},\cdots,e^{j\phi_{M_{\text{z}}}}\right]^{T}$.
@@ IV-A Joint Optimization
$\text{rank}(\boldsymbol{W})=$ $1$ (19)
@@
[…]
which is then iteratively solved using the SCA method[4]. Specifically, in the $-n\text{-}$th iteration, $\lVert\boldsymbol{W}\rVert_{2}$ is approximated by its first-order Taylor expansion at a given local point $\boldsymbol{W}^{n}$, i.e., $\lVert\boldsymbol{W}^{n}\rVert_{2}+\text{Re}\big{(}\text{Tr}\big{(}\left(\partial_{\boldsymbol{W}^{n}}\lVert\boldsymbol{W}\rVert_{2}\right)(\boldsymbol{W}-\boldsymbol{W}^{n})\big{)}\big{)}$, where the sub-gradient of $\lVert\boldsymbol{W}\rVert_{2}$ at $\boldsymbol{W}^{n}$ can be calculated as $\partial_{\boldsymbol{W}^{n}}\lVert\boldsymbol{W}\rVert_{2}=\boldsymbol{v}\boldsymbol{v}^{H}$, with $\boldsymbol{v}$ being the dominant singular vector of $\boldsymbol{W}^{n}$. Then, each iteration solves a convex subproblem, i.e.,
@@ IV-B Alternating Optimization (AO)
11:   Obtain $\boldsymbol{W}_{\text{z}}$ via similar procedures like steps 4$\sim$8.

### v2

@@ II Phase-Tuning Mechanism and Design of QS-IRS
Figure 1: (a) RA manufactured by TMYTEK comprising $\times51$ static elements with different patch shapes/sizes[17]; (b) RA with hybrid polarization-phase tuning based on element mirroring/rotation[18]; (c) Electromechanical IRS with rotatable patches[19]; (d) Electromechanical IRS with movable element ground[20].
@@ III System Model
Consider far-field communications between the BS and a target area on the ground with severely blocked direct links. We focus on practical scenarios where a QS-IRS is appropriately deployed [^3] The case with multiple distributed IRSs[22] is left for future investigation. such that there exists LoS paths from the BS to the target area via the QS-IRS, as in [2],[4]. [^4] The non-LoS scenario with non-negligible direct links will be considered in our future work, whereby statistical analysis of link power and outage probability under (e.g., Rician) fading channel can be incorporated in our system model and optimization framework. Consider one omnidirectional antenna at the receiver, and one (synthesized) directional antenna (beam) at the BS with a peak power gain of $G_{\text{t}}$ pointing towards the QS-IRS. [^5] Our proposed method can be fully protocol-transparent as in [6], under given BS antenna configuration. Assume that $M\triangleq M_{\text{y}}\times M_{\text{z}}$ IRS elements are arranged as a uniform planar array (UPA) with $M_{\text{y}}$ and $M_{\text{z}}$ elements along the $y\text{-}$ and $z\text{-}$axis, respectively, and with the first element at the origin, as shown in Fig. 2. Consider the IRS ERP model [10] given by $GF(\Pi)$, where $G$ denotes the peak power gain and $F(\Pi)$ denotes the normalized power pattern from/to the azimuth and elevation angles $\Pi\triangleq(\varphi,\theta)$, i.e.,
@@ III-A Array Steering Vector
Denote $\boldsymbol{u}_{\text{i}}\triangleq[\cos\varphi_{\text{i}}\sin\theta_{\text{i}},\sin\varphi_{\text{i}}\sin\theta_{\text{i}},\cos\theta_{\text{i}}]^{T}$ as the unit direction vector of the BS position w.r.t. the origin, with $\Pi_{\text{i}}\triangleq(\varphi_{\text{i}},\theta_{\text{i}})$ denoting the corresponding (horizontal, vertical) angle. The components of the steering vector incident on the QS-IRS along the $y\text{-}$ and $z\text{-}$axis can be expressed as
@@
where $\boldsymbol{r}_{m_{\text{y}}}\triangleq[0,(m_{\text{y}}-1)d_{\text{y}},0]^{T}$ and $\boldsymbol{r}_{m_{\text{z}}}\triangleq[0,0,(m_{\text{z}}-1)d_{\text{y}}]^{T}$ denote the position vectors of the $(m_{\text{y}},m_{\text{z}})$-th element along the $y\text{-}$ and $z\text{-}$axis, with $d_{\text{y}}$ and $d_{\text{z}}$ being the element spacing; $c$ denotes the speed of light and $f_{\text{c}}$ denotes the carrier frequency. Then, the incident steering vector is given by $\boldsymbol{a}_{\text{i}}(\Pi_{\text{i}})\triangleq\boldsymbol{a}_{\text{i}\text{y}}(\Pi_{\text{i}})\otimes\boldsymbol{a}_{\text{i}\text{z}}(\Pi_{\text{i}})$. Similarly, the components of the reflect steering vector are given by
@@ III-B Channel Power Gain
where (11) is due to (III-A) and the decomposition of $\boldsymbol{w}=\boldsymbol{w}_{\text{y}}\otimes\boldsymbol{w}_{\text{z}}$, with $y\text{-}$ component $\boldsymbol{w}_{\text{y}}\triangleq\left[e^{j\phi_{1}},\cdots,e^{j\phi_{M_{\text{y}}}}\right]^{T}$ and $z\text{-}$ component $\boldsymbol{w}_{\text{z}}\triangleq\left[e^{j\phi_{1}},\cdots,e^{j\phi_{M_{\text{z}}}}\right]^{T}$.
@@ IV-A Joint Optimization
$\text{rank}(\boldsymbol{W})=$ $1.$ (19)
@@
$\max_{\boldsymbol{W},\rho}\,\,10\log_{10}\rho-\sigma(\lVert\boldsymbol{W}\rVert_{*}-\lVert\boldsymbol{W}\rVert_{2}),$ (21)
which is then iteratively solved using the SCA method[4]. Specifically, in the $n\text{-}$th iteration, $\lVert\boldsymbol{W}\rVert_{2}$ is approximated by its first-order Taylor expansion at a given local point $\boldsymbol{W}^{n}$, i.e., $\lVert\boldsymbol{W}^{n}\rVert_{2}+\text{Re}\big{(}\text{Tr}\big{(}\left(\partial_{\boldsymbol{W}^{n}}\lVert\boldsymbol{W}\rVert_{2}\right)(\boldsymbol{W}-\boldsymbol{W}^{n})\big{)}\big{)}$, where the sub-gradient of $\lVert\boldsymbol{W}\rVert_{2}$ at $\boldsymbol{W}^{n}$ can be calculated as $\partial_{\boldsymbol{W}^{n}}\lVert\boldsymbol{W}\rVert_{2}=\boldsymbol{v}\boldsymbol{v}^{H}$, with $\boldsymbol{v}$ being the dominant singular vector of $\boldsymbol{W}^{n}$. Then, each iteration solves a convex subproblem, i.e.,
@@ IV-B Alternating Optimization (AO)
[…]